OpenStudy (anonymous):

0 divided by 6 7/12 [show steps please]

OpenStudy (anonymous):

0

OpenStudy (saifoo.khan):

\[\frac{0}{anything} = zero\]

OpenStudy (anonymous):

^what saifoo said.

OpenStudy (anonymous):

anything/0 = EXPLOSIONS EVERYWHERE

OpenStudy (saifoo.khan):

Even on Mars.

OpenStudy (saifoo.khan):

Even on Saturn.

Directrix (directrix):

0 divided by( 6 7/12) = 0 / 79/12 = 0 * 12/70 = 0

OpenStudy (anonymous):

Alright, this problem is in my Supplementary Problems section and the answer I'm getting from them is 2. That's why I asked. So confused...

OpenStudy (anonymous):

@funinabox I'm pretty sure it's black holes not explosions...

OpenStudy (saifoo.khan):

And on Sun, it's Always \[\frac{hydrogen}{zero} = Expositions\]

OpenStudy (anonymous):

oooohhhh....

OpenStudy (anonymous):

@saifoo.khan You are wrong because 0/0 =/= 0

OpenStudy (anonymous):

If you are going to state a math rule at least explain it fully so it doesn't lead to confusion.

OpenStudy (saifoo.khan):

\[\frac00 \to \frac{expositions}{expositions} \to \text{indeterminate Explosions}\]

Directrix (directrix):

Look at the problem again to ensure that it is written correctly. I viewed 6 7/12 as a mixed number. Is that correct? I will work with you after these people finish acting the fool. But, please check the problem

OpenStudy (anonymous):

0 divided by 6 7/12 (mixed number, yes) ..the book is saying 2. Ugh Lol.. I'm like HOW. I've been doing it and I always get zero. So confusing. Maybe there's an error in the book?

OpenStudy (saifoo.khan):

That dosnt matter, @LynnMichael , you can answer this question by just looking it at a glance. like i said, \[\frac{0}{anything} = Zero\]No need to go into hardships.

OpenStudy (anonymous):

I know @saifoo.khan, I'm just wondering why the book is saying something else. Thank you though.

Directrix (directrix):

@LynnMichael Please check the problem to see if you wrote it correctly. Thanks.

Directrix (directrix):

0 divided by 6 7/12 (mixed number, yes)

Directrix (directrix):

0 /( 6 and 7/12) = 0 / [72/12 + 7 / 12] =

Directrix (directrix):

0 / (79/12) = 0